What is the correct value for operating expenses? i got 317200 only by calculating bakcwards from every other value. Also are the rest of my values correct? i'm not sure if investing and financing are correctHere is a statement of financial position for Big Bang Balloons Pty Ltd for 2015 and comparative figures for 2014 Additional
Answer #1
Computation of cash paid for Operating expenses
Total operating expenses (Including depreciation):
$ $
Total expenses 1047600
Less:
Cost of sales expense 633600
Interest expense 14000
Tax expense 54000 701600
Total operating expenses 346000
Depreciation=Increase in accumulated depreciation=79200-50400=$ 28800
Cash paid for operating expenses=Total operating expenses-Depreciation=346000-28800=$ 317200
Rest of your values are correct